# Spectraformer: A Unified Random Feature Framework for Transformer

### v3

@@ A.2.1 Alternative Kernelized Attentions
A.2.2 Kernel Approximation Beyond Random Features
Kernel approximation is presented in Section 2.3 through the use of random features. However, there has been other kernel approximation techniques outside of random features. These include greedy basis selection (Smola & Schökopf, 2000), divide and conquer (Hsieh et al., 2014; Zhang et al., 2013), and nyström methods (Williams & Seeger, 2000) (which led to the development of Skyformer (Chen et al., 2021) and Nyströmformer (Xiong et al., 2021)).
Figure 2: Spectraformer weight matrices. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work. Please refer to Appendix A.3 for references.
inline]MAYBE: Table here like table 2 comparison of different kernel space and time complexities combined with table 3 comparison of convergence rates and required random features for kernel approximation error in (Liu et al., 2022)
A.3 Technical Details of Weight Approximators
Spectraformer weight matrices are shown in Figure 2. Weight matrices are either approximators or learners. There are three approaches to approximators: Monte Carlo sampling, Unit-cube sampling, and quadrature. We have discussed these approaches in detail in Section 3.1. Here we provide a more specific classification schemes for each method.
  • Monte Carlo sampling-based weight matrices are improvements over Base and subdivided into two types: structural and geometric methods. Structural methods decompose the Base random matrices $P$ into smaller matrices which reduce time and/ or space complexity. Structural matrices includes FastFoodF, SCRF and P-Model. Among these, we have explored the learner variant FastFoodL of FastFood. Geometric methods enforce certain geometrical couplings along some dimensions in the construction of random matrices to reduce the approximation error. Geometric methods include: ROM (ORF, SORF) and SimRF. Among these, we have explored the main ROM variants - ORF and SORF.
  • Unit-cube sampling-based weight matrices reformulate the Bochner integral in Equation 9 as Equation A.3 with uniform and independent points. This approach includes QMC, MM, and SSF. We have not covered SSF in our experiments.
  • Quadrature-based weight matrices estimates the intractable integral using non-uniform weights with deterministic rules. There are two approaches: SGQ and SSR. We have not covered SSR in our experiments.
inline]update the equations so that we use the same dimension for the number of components and so on inline]Generative should be DGM, add Base to the graph
We now explore these weight matrices in the order specified above.
Base (Rahimi & Recht, 2007) is the direct sampling from a distribution $\bm{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
$W_{Base}=\frac{1}{\sigma}\bm{P}$ (16)
$W_{FastFood_{L}}=\frac{1}{\sigma}\bm{SHG\Pi HB}$ (17)
[…]
SCRF (Signed Circulant Matrix Projection) (Feng et al., 2015) maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF. It does this by making use of circulant matrices.
$\bm{W}_{SCRF}=[\bm{P}^{(1)},\bm{P}^{(2)},...,\bm{P}^{(t)}],t=d|D$ (18)
where $\bm{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\bm{C}^{(i)_{[d]}}$ of $\bm{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}/\sigma^{2})$. The circulant matrix via Discrete Fourier Transform can be defined as:
$\begin{array}[]{cc}\bm{C}=\frac{1}{m}\bm{F}^{*}diag(\bm{Fc})\bm{F}\\ \bm{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
P-Model (Choromanski & Sindhwani, 2016) generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
$\bm{W}_{P_{model}}=[\bm{g^{T}P_{1}},\bm{g^{T}P_{2}},...,\bm{g^{T}P_{s}}]^{T}$ (20)
where $\bm{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\bm{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model. Semi-Gaussian structured matrices including Toeplitz, Hankel, and other Toeplitz-like matrices are an open area of research for random matrices.
ORF (Orthogonal RF) (Yu et al., 2016) decreases the approximation error significantly compared to Base. This is done via replacing $G$ with a properly scaled random orthogonal matrix. However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
$\bm{W}_{ORF}=\frac{1}{\sigma}\bm{SQ}$ (21)
SORF (Structured ORF) (Yu et al., 2016) decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods. SORF is unbiased with large $d$. We replace $\bm{S}$ with $\sqrt{d}$ and $\bm{Q}$ with structured matrix $\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
$\bm{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$ (22)
ROM (Random Orthogonal Embeddings) (Choromanski et al., 2017) generalizes SORF ($t=3$) to:
$\bm{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\bm{SD}_{i}$ (23)
where $\bm{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\bm{H}$ is a representative of, and $\bm{D}$ being independent diagonal matrices. When $\bm{S}=\bm{H}$ and $t=3$, we obtain SORF used in our experiments. Another ROM variant is S-Hybrid, which we do not experiment with, given as
$\bm{W}_{S_{Hybrid}}=\bm{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\bm{SD}_{i}^{(\mathcal{R})}$ (24)
where $\bm{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d. $Unif(\mathcal{S}^{1})$ on the unit circle of $\mathbb{C}$. When being used in the component function, instead of calculating $\phi(\bm{x})^{T}\phi(\bm{y})$ in Equation 10. We model the equation as $Real(\overline{\phi(\bm{x})}^{T}\phi(\bm{y}))$.
SimRF (Simplex RF) (Reid et al., 2023) is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods. This is an alternative to the ROM-based geometric methods. Two variants of SimRF are: weight-independent and weight-dependent. SimRF is defined as:
$\bm{W}_{simp}=\bm{DSR}$ (25)
where $\bm{D}\in\mathbb{R}^{d*d}=diag(\bm{\omega}_{i}),\bm{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\bm{S}\in\mathbb{R}^{d*d}$ with row $\bm{s}_{i}$:
$\bm{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\bm{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
being unit vectors which are manifestly normalized and subtend obtuse angles. SimRF variants are defined based on how $\bm{\omega}_{i}$ is constructed. For SimRFindep, $\bm{\omega}_{i}\sim\chi_{d}$. For Simdep, we permit the random vector direction $\{\hat{\bm{\omega}}_{i}\}$ to be correlated with norms $\{\bm{\omega}_{i}\}$:
$\bm{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\bm{\omega}_{i},i=1,...,d$ (27)
QMC (Quasi-Monte Carlo) (Avron et al., 2016) evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo. Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect. To calculate QMC, we first assume that $p(\bm{x})=\prod^{d}_{j=1}p_{j}(\bm{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function. Then we define:
$\Phi^{-1}(\bm{t})=\left(\Phi_{1}^{-1}\left(\bm{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\bm{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$. We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $[0,1]^{d}$ as
The weight matrix then can be defined as:
MM (Moment Matching) (Shen et al., 2017; Liu et al., 2022) improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features. This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
where $\widetilde{\Phi}^{-1}(\bm{t}_{i})=\tilde{\bm{A}}^{-1}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\bm{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\bm{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\bm{A}}$ satisfying $\tilde{\bm{A}}\tilde{\bm{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$.
SSF (Spherical Structured Features) (Lyu, 2017) improves over QMC by including rotation-invariant kernel as well as in terms of complexity. Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\bm{x}\in\mathbb{R}^{d+1}|||\bm{x}||_{2}=1\}$. Specifically, we construct $\{\bm{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\bm{V}:=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
$\bm{W}_{\text{SSF}}=[\Phi^{-1}(t)\bm{v}_{1},\Phi^{-1}(t)\bm{v}_{2},\ldots,\Phi^{-1}(t)\bm{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
SGQ (Sparse Grid Quadrature) (Dao et al., 2017) evolves from GQ. GQ (Gaussian Quadrature) (Dao et al., 2017) assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule. However the total number of points $s$ scale exponentially with the dimensions. However GQ suffers form the curse of dimensionality. This is alleviated using Smolyak rule resulting in SGQ. Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
SSR (Spherical-Radial Rules) (Munkhoeva et al., 2018) transforms Equation 10 into a double integral over a hyper-sphere and the real line. This leads to the following approximation:
$\bm{W}_{\text{SSR}}=\bm{\vartheta}\otimes\begin{bmatrix}(\bm{QV})^{\top}\\ -(\bm{Q}\bm{V})^{\top}\end{bmatrix}\in\mathbb{R}^{(2d+1)\times d},$
where $\bm{\vartheta}=[\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{s}]$, $\bm{V}=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{d+1}]$, $\vartheta\sim\chi(d+2)$ and $\{\bm{v}_{i}\}_{i=1}^{d+1}$ being vertices of a unit regular d-simplex randomly rotated by $\bm{Q}$ (a random orthogonal matrix).
A.4 Technical Details of Weight Learners
We give a non-mathematical definition of weight matrix learner in Section 3.1: weight matrix learner is any function which parameterize distributions. This allows for the discovery and future experimentation with other potential weight matrix learners in Spectraformer and random feature method in general. Weight matrix learner comes from a long line of work in kernel methods, namely, kernel learning via random features. Typically, a component function $f$ is chosen and a weight matrix $\bm{W}$ is produced via kernel approximation, providing us with the feature map $\phi$. $\phi$ can then be combined with input $\bm{x}$ and weight $\bm{v}$ to fit the learning objective. Kernel learning, however, imposes an additional objective: learning the weight matrix $\bm{W}$.
The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods (Liu et al., 2022). Two-stage methods (e.g., Yu et al. (2015); Wilson & Adams (2013); Bullins et al. (2018)) solves the dual objective separately: $\bm{W}$ is typically first learned via a kernel alignment scheme (Wang et al., 2015), then we solve for $\bm{v}$. One-stage methods (e.g., Chowdhury et al. (2022); Tompkins et al. (2019)), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\bm{W}$ is parameterized then $\bm{v}$ is solved typically. It is not plausible to apply kernel alignment (Wang et al., 2015) on Transformer architectures. Therefore, only one-stage methods can be considered in the context of weight matrix learner. Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting (Yang et al., 2015), then adapted into the Transformer setting (Chowdhury et al., 2022). DGM (Deep Generative Model) has also been effective in modeling distributions (Kingma & Welling, 2013). It too has been adapted by Chowdhury et al. (2022) for the Transformer setting. Due to the superior performance of FastFoodL with this technique, we have only experimented with this technique. We now explore them.
FastFoodL (see Equation 17) is proposed to be learnable by Yang et al. (2015) with two variations: FSARD (scaling matrix $\bm{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\bm{G}$ and $\bm{B}$). It is further adapted by Chowdhury et al. (2022) to make $\bm{S}$ or $\bm{S},\bm{G},\bm{B}$ learnable parameters. In our experiments, we maintain this set up.
GMM (Gaussian Mixture Model) (Wilson & Adams, 2013; Oliva et al., 2016) is a known universal approximator of probability distribution. Specifically, we can consider $\bm{W}_{GMM}$ being composed of $s$ components, with each component $c$ (with $C$ total components) being a Gaussian with mean vector $\mu_{c}$ and covariance matrix $\Sigma_{c}$ and component weight $\pi_{c}$. $p(\bm{\omega})$ can then be approximated as $p(\bm{\omega})=\sum_{c=1}^{C}\pi_{c}\mathcal{N}(\mu_{c},\Sigma_{c})$ (Chowdhury et al., 2022). Assuming $\pi_{c}=\frac{1}{C}$, the weight matrix with learnable weights $\Sigma$ and $\mu$ can be parameterized as:
$\bm{W}_{GMM}=[\Sigma_{1}n+\mu_{1},...,\Sigma_{C}n+\mu_{C}]^{T},n\sim\mathcal{N}(\bm{0},\bm{I})$ (29)
Another formulation of GMM is proposed in Yang et al. (2015) in combination with FastFood, albeit it extends beyond constructing the weight matrix.
DGM (Deep Generative Model) (Kingma & Welling, 2013) allows for a generic distribution modeling using neural network. DGM is initially proposed to make differentiating sampling processes possible via the reparametrization trick, where the generator $g$ creates samples from the target distribution $\bm{\omega}_{m}=g(\bm{n}_{m})$ via transforming from a noise distribution $\bm{n}_{m}\sim p_{0}(.)$.
In addition to the techniques discussed above, there are also some other prominent one-stage methods (e.g., INN, BBQ). INN (Invertible Neural Network) (Xu et al., 2021) is another universal approximator of probability distribution, consisting of invertible operations transforming samples from a known distribution to a different more complex distribution. BBQ (Black Box Quantile) (Tompkins et al., 2019) follows the QMC scheme and models $\Phi^{-1}$ using parameterized quantile function.
A.5 Technical Details of Component Functions
Figure 3: Spectraformer component functions. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work. Please refer to Appendix A.5 for references.
Component functions discussed in Section 3.2 all belong to CDERF. However, there are other component functions including DIRF and HRF (see Figure 3) which we do not explore in our experiments. We now investigate the theoretical details of these component functions, the CDERF family is introduced in their order of discovery.
CDERF (Complex dense exponential random feature) is given in Equation 30, where $\bm{\omega},\bm{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of $d*d$ complex symmetric matrices. The parameters and constraints of different CDERF functions are specified in Table 7:
$f_{DE}^{(k)}(\bm{\omega},\bm{x})=Real(D\exp(\bm{\omega}^{T}\bm{A}\bm{\omega}+\bm{\omega}^{T}\bm{B}^{(k)}\bm{x}+\bm{x}^{T}\bm{C}^{(k)}\bm{x}))$ (30)
Table 7: CDERF component functions
$\bm{A}$ $\bm{B}^{(1)}$ $\bm{B}^{(2)}$ $\bm{C}^{(1)}$ $\bm{C}^{(2)}$ $D$ $\bm{x}$ $\bm{y}$
[…]
TrigRF (Rahimi & Recht, 2007) $0\bm{I}_{d}$ $i\bm{I}_{d}$ $-i\bm{I}_{d}$ $0\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
PosRF (Choromanski et al., 2021) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
OPRF (Likhosherstov et al., 2022) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $-1\bm{I}_{d}$ $1$ $\bm{x}$ $\bm{y}$
TrigRF (Trigonometric RF) (Rahimi & Recht, 2007) is the base RFF implementation.
$\begin{array}[]{c}f_{TrigRF}^{(1)}(\bm{\omega},\bm{x})=\sqrt{2}\cos(\bm{\omega}^{T}\bm{x}+\bm{b})\\ f_{TrigRF}^{(2)}(\bm{\omega},\bm{y})=\sqrt{2}\sin(\bm{\omega}^{T}\bm{y}+\bm{b})\\ \bm{b}\sim Uniform(0,2\pi)\end{array}$ (31)
PosRF (Positive RF) (Choromanski et al., 2021) fixes the problem of TrigRF by enforcing positive component function output in the softmax. The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0. PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 32) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 33).
$\begin{array}[]{c}f_{PosRF_{B}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-\frac{||\bm{x}||^{2}}{2})\end{array}$ (32)
$\begin{array}[]{ll}f_{1,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\\ f_{2,PosRF_{Hyp}}(\bm{\omega},\bm{x})=\exp(-\bm{\omega}^{T}\bm{x}-||\bm{x}||^{2})\end{array}$ (33)
We only use PosRF-B in our experiments.
GERF (Generalized exponential RF) (Likhosherstov et al., 2022) generalizes both TrigRF and PosRF with Equation 34.
$\begin{array}[]{c}f_{GERF}^{(1)}(\bm{\omega},\bm{x})=D\exp(A||\bm{\omega}||^{2}+B\bm{\omega}^{T}\bm{x}+\bm{C}||\bm{x}||^{2})\\ f_{GERF}^{(2)}(\bm{\omega},\bm{y})=D\exp(\bm{A}||\bm{\omega}||^{2}+sB\bm{\omega}^{T}\bm{y}+\bm{C}||\bm{y}||^{2})\\ \\ Re(1-4A)>0,\bm{B}=\sqrt{s(1-4\bm{A})}\\ \bm{C}=-(s+1)/2,D=(\sqrt[4]{1-4\bm{A}})^{d}\\ \bm{A}\in\mathbb{C},s\in\{-1,+1\}\end{array}$ (34)
where $\sqrt{.}$ and .$n]{.}$ denoting a principal root with a complex argument
OPRF (Optimized positive RF) (Likhosherstov et al., 2022) is the solution to the minimization of the variance of GERF. Specifically it is defined as Equation 34 with $s=+1$, $||\bm{x}+\bm{y}||^{2}>0$, and $\bm{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
$\begin{array}[]{cc}\bm{A}=(1-1/p^{*})/8\\ p^{*}=(\sqrt{(2||\bm{x}+\bm{y}||^{2}+d)^{2}+8d||\bm{x}+\bm{y}||^{2}}-2||\bm{x}+\bm{y}||^{2}-d)\\ /(4||\bm{x}+\bm{y}||^{2})\end{array}$ (35)
DERF (Dense-exponential random features) (Likhosherstov et al., 2023) extends GERF and replace $\bm{A},\bm{B},\bm{C}$ with dense matrices. DERF is CDERF when $\bm{B},C$ are in the real instead of the complex plane. Minimizing the variance of DERF leads to two approaches: ADERF and SDERF. However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries. Therefore SADERF is proposed.
SADERF (simplified ADERF) (Likhosherstov et al., 2023) is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
$\begin{array}[]{c}f_{SADE}^{(1)}(\bm{\omega},\bm{x})=f_{GE}^{(1)}(\bm{\omega},\Psi\bm{x}),f_{SADE}^{(2)}(\bm{\omega},\bm{y})=f_{GE}^{(2)}(\bm{\omega},\Psi^{-1}\bm{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\bm{y}_{l}^{(i)})^{2}/\sum_{i}(\bm{x}_{l}^{(i)})^{2})^{1/4}\end{array}$ (36)
DIRF (Discretely-induced random features) (Likhosherstov et al., 2022) are based on the assumption of $p(\bm{\omega})$ being a discrete distribution with $\bm{\omega}_{1},...,\bm{\omega}_{d}$ being i.i.d., $\mathbb{P}(\bm{\omega}_{l}=k)=p_{k}$, $\sum_{k=0}^{\infty}p_{k}=1$, $p_{k}>0,k\in\{0\}\cup\mathbb{N}$. By making use of the Taylor series, we can define DIRF as:
$f_{DI}(\bm{\omega},\bm{x})=\exp(-\frac{||\bm{x}||^{2}}{2})\prod_{l=1}^{d}\bm{x}_{i}^{\bm{\omega}_{l}}(\bm{\omega}_{l}!)^{\frac{-1}{2}}p_{\bm{\omega}_{l}}^{\frac{-1}{2}}$ (37)
DIRF has two variations: PoisRF (Poisson RF) which requires $p(.)$ to be the Poisson distribution and GeomRF (Geometric RF) which requires $p(.)$ to be the geometric distribution. Since the many weight matrices introduced in Section 3.1 do not readily approximate these distributions, we decide to leave DIRF out of the paper.
HRF (Hybrid RF) (Choromanski et al., 2022) combines multiple base estimators, similar to multiple kernel learning methods (Bach et al., 2004), in order to provide the most accurate approximation in regions of interest. Choromanski et al. on solving softmax approximation using HRF. HRF is constructed using a weighted combination $\epsilon$ of $p+1$ base estimators $\hat{SM}^{k}(\bm{x},\bm{y})$ and $\lambda$-coefficients, $\Lambda$ $p$ binary estimators $\lambda^{k}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow[0,1]$, both constructed independently:
$\hat{SM}^{\epsilon,\Lambda}=\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y})\hat{SM}^{k}(\bm{x},\bm{y})+(1-\sum_{k=1}^{p}\hat{\lambda}^{k}(\bm{x},\bm{y}))\hat{SM}^{p+1}(\bm{x},\bm{y})$ (38)
Since HRF requires the base estimators $SM(\bm{x},\bm{y})$ to be unbiased, the easiest solution is the complex exponential $\hat{SM}^{cexp}_{m}$ which can be derived directly from the expectation of the softmax as follows:
$\begin{array}[]{cc}\hat{SM}^{cexp}_{m}(\bm{x},\bm{y})=\psi^{m}_{A}(\bm{x})^{T}\psi^{m}_{(\bm{A}^{T})^{-1}}(\bm{y})\\ \psi^{m}_{\bm{M}}(u)=\frac{1}{\sqrt{m}}\exp(-\frac{(\bm{Mu})^{2}}{2})\\ (\exp(\bm{\omega}^{T}_{i}\bm{Mu}),...,\exp(\bm{\omega}_{m}^{T}\bm{Mu}))^{T},\bm{\omega}_{i}\sim\mathcal{N}(\bm{0},\bm{I}_{d})\end{array}$ (39)
To construct $\lambda$-coefficients, three methods are proposed: simultaneous accurate approximation of both small and large values, Gaussian Lambda Coefficients, and adaption to data admitting clustering structure.
A.6 Experimental Details in Spectraformer
Table 8: Hyper-parameters for Spectraformer, following the setting from Chen et al. (2021) due to limited computational power. $L$ stands for ListOps, $T$ for Text, and $R$ for Retrieval. The seed that can be used to reproduce our experiments is 42.
L T R
Embedding dim. 64 64 64
Transformer dim 64 64 64
Hidden dim 128 128 128
Head dim 32 32 32
Num. heads 2 2 2
Num. layers 2 2 2
Vocabulary size 32 512 512
Sequence length 2000 4000 4000
Dropout rate 0.1 0.1 0.1
Att. dropout rate 0.1 0.1 0.1

### v5

@@ 5.1. Alternative Kernelized Attentions
5.2. Kernel Approximation Beyond Random Features
Kernel approximation is presented in Section 2.3 through the use of random features based on Bochner’s theorem limiting to only symmetric, positive-definite (PD) kernels. While Spectraformer focuses on approximating the symmetric softmax kernel, other research has extended the RFF paradigm to handle other classes of functions.
Notably, the AsK-RFFs (Asymmetric Kernel RFFs) (He et al., 2024) method generalizes RFFs to handle asymmetric kernels. The key theoretical innovation is the use of a complex measure expanding on the Bochner theorem, which can be decomposed into four finite positive measures. This allows AsK-RFFs to create two distinct feature mappings that can approximate an asymmetric kernel. A different class of kernel functions are the Generalized Zonal Kernels (GZKs) which are dot-product kernels on the unit sphere, which is different from shift-invariant kernels which are the focal point of this paper. Nevertheless, the Gaussian is a part of this class. Random Gegenbauer Features (Han et al., 2022) are designed to approximate the GZKs. This is done by performing a low-rank approximation via sampling the GZK’s feature map, defined by an infinite series of Gegenbauer polynomials instead of using the entire feature map.
In addition, there has been other kernel approximation techniques outside of random features. These include greedy basis selection (Smola and Schökopf, 2000), divide and conquer (Hsieh et al., 2014; Zhang et al., 2013), nyström methods (Williams and Seeger, 2000) (which led to the development of Skyformer (Chen et al., 2021) and Nyströmformer (Xiong et al., 2021)). Nystrom methods are a type of low rank attention of which Low-Rank Transformer (Winata et al., 2020) is a part of. There is also sparse attention (Big Bird (Zaheer et al., 2020) which combines windowed, global and random attention, Reformer (Kitaev et al., 2020) which uses locality-sensitive hashing, Longformer (Beltagy et al., 2020) which combines band attention and global attention on special tokens like [CLS], ETC (Encoding Long and Structured Inputs) (Ainslie et al., 2020), Routing Transformer (Roy et al., 2021), Star Transformer (Guo et al., 2019)), time-series inspired methods (Informer (Zhou et al., 2021)), or Linformer (Wang et al., 2020). The inclusion of many of these models in our LRA benchmark comparison (Table 1) provides a broader comparison for evaluating the performance of Spectraformer against other efficient Transformers beyond random features. For a more comprehensive overview, please see the survey in (Elouargui et al., 2023).
An alternative to using random features is directly specifying $\phi(x)$ without approximating the softmax via random features. cosFormer (Qin et al., 2022) specifies $\phi(x)$ as LeakyReLU and ReLU.
6. Conclusion
We present Spectraformer, an open-source framework for approximating and learning the attention kernel in the Transformer. Our paper generalizes past works, and presents empirical findings on different component function and weight matrix combinations. We experiment with 18 combinations, of which 14 are novel. Spectraformer has produced variants that are competitive against a wide range of efficient Transformers, including those based on sparse attention and other approximation methods as shown in our expanded LRA benchmark. We have demonstrated that a novel combination discovered through our framework, OPRF-FastFoodL, performs on par with leading models like Nystromformer and BigBird, establishing a new SOTA among random feature-based methods and a strong contender in the overall efficient Transformer landscape. We suggest the use of weight matrices FastFoodL and QMC combined with component functions based on specific computational requirements (PosRF for less and OPRF for more computational requirement). Our work shows the validity of Spectraformer for studying and producing novel variants based on existing and future research in random features and distribution sampling.
7. Future work, limitations, and broader impact
Spectraformer provides an experimental framework for random features in efficient Transformer with kernelized attention. We are aware that many works cited do not provide comparable error approximation bounds. In addition, as (Yao et al., 2023) point out, most theoretical bounds are either highly conservative or involving unknown qualities, therefore making comparisons impossible. Therefore, we have validated our framework on extensive tasks, and provided comprehensive analyses and intuitive explanation. We welcome future work to generalize and build on Spectraformer. We have experimented with a significant number of combinations but not the exhaustive list in Sections 3.1 and 3.2. Future experiments can work with these additional techniques. In addition, providing a definitive theoretical reason for the success of specific method-task combinations, such as PosRF-QMC on ListOps, remains a significant open question that aligns with broader challenges in the kernel and random features literature. A rigorous theoretical investigation into these interactions is a valuable direction for future work.
Transformer-based architecture can also benefit from Spectraformer. Past random feature Transformer works (Likhosherstov et al., 2022, 2023) have experimented with conformer and vision Transformer with ViT. Jeevan and Sethi (2022) provide a good foundation to extend our work. Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification. It could also benefit from ablation studies like the comparison variance in (Chowdhury et al., 2022) or in (Yao et al., 2023). Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly (Weidinger et al., 2021).
Acknowledgments
The production of the code used in this paper was partially reliant on generative AI code assistant. All AI-generation has been validated by the authors. This research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. We would also like to acknowledge the ARC Centre of Excellence for Automated Decision Making and Society (CE200100005).
Appendix
Table 6. Hyper-parameters for all Spectraformer experiments on the Long Range Arena (LRA) benchmark tasks. The settings follow Chen et al. (2021) to ensure a fair comparison under limited computational power.
ListOps (L) Text (T) Retrieval (R) Image (I) Pathfinder (P)
Embedding dim. 64
Transformer dim 64
Hidden dim 128
Head dim 32
Num. heads 2
Num. layers 2
Vocabulary size 32 512 512 256 512
Sequence length 2000 4000 8000 1024 1024
Dropout rate 0.1
Att. dropout rate 0.1